Circular No. 8524

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

S/2005 S 1 C. C. Porco, CICLOPS, Space Science Institute, Boulder; and the

Cassini Imaging Science Team report the discovery of a new satellite of Saturn, designated S/2005 S 1, orbiting within the Keeler gap in Saturn’s outer A ring. (The object had been previously inferred from the presence of features observed on the outer edge of the Keeler gap; cf. Porco et al. 2005, Science 307, 1226, Fig. 10). The object was discovered in six images taken over 16 min on May 1 from a time-lapse sequence of 0.180-s narrow-angle-camera exposures that were targeted to the illuminated side of the outer edge of the A ring (with phase angle ∼ 33^{o} and image scale 6.9 km/pixel). S/2005 S 1 was subsequently found in 32 (7 km/pixel) low-phase images taken of the F ring on Apr. 13 (spanning 18 min) and again in two highresolution (3.54 km/pixel) low-phase images taken on May 2, when its 7-km disk was resolved. The satellite orbits Saturn every 0.594 day at a distance of 136500 km. The estimated geometric albedo is 0.5. The data are too coarse to yield any statistically significant orbital eccentricity or inclination.

V2361 CYGNI R. W. Russell, R. J. Rudy, and D. K. Lynch, Aerospace Corporation;

and W. Golisch, Infrared Telescope Facility (IRTF), report on a 0.8–2.5-μm spectrogram taken of V2361 Cyg (cf. IAUC 8483) with the IRTF (+ SpeX) on Mar. 6.66 UT. The calibrator was HD 192538 (an A0 star with mag V ∼ 6.5). The most prominent aspect of the spectrum was a strong, smooth continuum rising toward longer wavelengths across the entire spectral range — indicative of thermal emission from dust — and the continuum was wellfitted by a 970-K (± 10 K) blackbody. Also present in the spectrum were the following doubled emission lines (2600 km/s FWHM) with prominent red components: He I 1.0830-μm, H I 1.2818-μm, and O I at 0.8446, 1.1287, and 1.3165 μm. H I Paγ was present but very weak and blended with He I. The reddening, as determined from the O I lines, was E(B − V) ∼ 1.2 mag, a significant fraction of which may be local to the source.

COMET C/1998 U7 (SOHO) Another faint Kreutz sungrazer, stellar in appearance (cf. IAUC 8519):

Comet 1998 UT α_{2000} δ_{2000} Inst. F MPEC
C/1998 U7 Oct. 26.979 13^{h}53.^{m}4 −15^{o}29′ C3 XL 2005-H24

2005 May 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2005 CBAT  Daniel W. E. Green